OpenStudy (anonymous):

find the a so the limit is equal to e lim x -> inf ((x+a)/(x-a))^x find the a so the limit is equal to e lim x -> inf ((x+a)/(x-a))^x @Mathematics

OpenStudy (anonymous):

I'm pretty sure that the a should be 1/2 by guessing but how can I show it?

OpenStudy (zarkon):

it is not 1/2

OpenStudy (zarkon):

oh...sorry misread the problem...

OpenStudy (zarkon):

I was just taking the limit...got \[e^{2a}\]

OpenStudy (anonymous):

so if a is 1/2 then the limit would be e like I need right?

OpenStudy (zarkon):

I thought that is what you wanted...clearly 1/2 is the answer

OpenStudy (anonymous):

I guessed to get that I don't know how to solve the problem, I only have the solution

OpenStudy (anonymous):

I want to know how I can find that solution from the problem

OpenStudy (anonymous):

wait how did you evaluate the limit?

OpenStudy (zarkon):

are you allowed to use \[\lim_{x\to\infty}\left(1+\frac{r}{x}\right)^x=e^r\]

OpenStudy (anonymous):

is that an identity?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

I've never seen that before I think we're supposed to be using l'hopital's rule to solve this

OpenStudy (zarkon):

normally written \[\lim_{n\to\infty}\left(1+\frac{x}{n}\right)^n=e^x\]

OpenStudy (anonymous):

is there a way to solve it without using that? or can you show why that identity works?

OpenStudy (zarkon):

have you seen this \[\lim_{n\to\infty}\left(1+\frac{1}{n}\right)^n=e\] this is a definition of e

OpenStudy (anonymous):

yeah I have seen that before

OpenStudy (zarkon):

I can derive the identity from there ,but if you are supposed to use l'hospitals rule we can do that

OpenStudy (anonymous):

well that's what we learned this week so could we please do it that way?

OpenStudy (anonymous):

I think I've made some progress I have\[\lim_{x \rightarrow \infty} e ^{xln(x+a/x-a)}\]

OpenStudy (zarkon):

write \[\left(\frac{x+a}{x-a}\right)^x\] \[e^{x\ln\left(\frac{x+a}{x-a}\right)}\]

OpenStudy (zarkon):

good

OpenStudy (anonymous):

then was thinking that the top needed to be 1 as x->inf

OpenStudy (anonymous):

so I focused on that

OpenStudy (anonymous):

but the limit of that is indeterminate

OpenStudy (anonymous):

but not in the right way to use l'hopital's rule I don't think

OpenStudy (zarkon):

\[e^{x\ln\left(\frac{x+a}{x-a}\right)}\] \[=e^{\frac{\ln\left(\frac{x+a}{x-a}\right)}{1/x}}\]

OpenStudy (zarkon):

compute \[\lim_{x\to\infty}=\frac{\ln\left(\frac{x+a}{x-a}\right)}{1/x}\]

OpenStudy (anonymous):

and that puts in 0/0 form right?

OpenStudy (zarkon):

if you do it right you will get 2a .... and so the final answer will be \[e^{2a}\]

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

can you show me how to take the derivative of that?

OpenStudy (anonymous):

things don't seem to be getting any simpler the quotient rule means that I keep the ln(x+a/x-a)

OpenStudy (anonymous):

then I just have a big mess from the ln(x+a/x-a) which also uses the quotient rule

OpenStudy (zarkon):

write \[\frac{x+a}{x-a}\] as \[\frac{x+a}{x-a}=\frac{x-a+2a}{x-a}=1+\frac{2a}{x-a}\]

OpenStudy (anonymous):

oh good idea I didn't think of doing that

OpenStudy (zarkon):

\[\frac{d}{dx}\ln\left(1+\frac{2a}{x-a}\right)\] \[=\frac{1}{1+\frac{2a}{x-a}}\frac{-2a}{(x-a)^2}\]

OpenStudy (zarkon):

\[\frac{d}{dx}\frac{1}{x}=\frac{-1}{x^2}\]

OpenStudy (zarkon):

put together...

OpenStudy (zarkon):

\[\frac{\frac{1}{1+\frac{2a}{x-a}}\frac{-2a}{(x-a)^2}}{\frac{-1}{x^2}}\]

OpenStudy (zarkon):

\[=\frac{1}{1+\frac{2a}{x-a}}\frac{-2a}{(x-a)^2}(-x^2)\] \[=\frac{1}{1+\frac{2a}{x-a}}\times2a\times\frac{x^2}{(x-a)^2}\]

OpenStudy (zarkon):

now take limit

myininaya (myininaya):

good job in turning that warlock into a princess

OpenStudy (zarkon):

lol

OpenStudy (anonymous):

really that was amazing

myininaya (myininaya):

more and more practice you will get better at being manipulative

OpenStudy (anonymous):

but for now it looks like magic lol

OpenStudy (zarkon):

you understand what I typed?

myininaya (myininaya):

well turning a warlock into a princess is just a trick its not really magic

OpenStudy (anonymous):

I do have one question actually, how come you didn't need to use the quotient rule when you put it together?

OpenStudy (anonymous):

you just took the derivative of the top and put it over the derivative of the bottom

OpenStudy (zarkon):

That is L'Hospitals rule

OpenStudy (anonymous):

ohh yeah, no wonder I was so messed up haha

OpenStudy (zarkon):

;)

OpenStudy (anonymous):

it all makes sense now! Thank you so much!

OpenStudy (zarkon):

good

myininaya (myininaya):

you can ask zarkon any math question hes a math god

OpenStudy (anonymous):

he really is lol

OpenStudy (zarkon):

Demigod

OpenStudy (zarkon):

there are still things I don't know

myininaya (myininaya):

not half maybe three fourths

OpenStudy (zarkon):

lol..I'll take that

myininaya (myininaya):

\[\frac{3}{4}-god\]

myininaya (myininaya):

oops that looks like a minus sign

OpenStudy (zarkon):

it does

myininaya (myininaya):

it was suppose to be a dash

myininaya (myininaya):

you know anything about Diffie-Hellman tuples?

OpenStudy (zarkon):

I'm not familiar with Diffie-Hellman (I know what a tuple is ) :)

myininaya (myininaya):

there is alittle more to it but i know it has something to do with cyclic groups

OpenStudy (zarkon):

ok...I know what a cyclic group is

OpenStudy (zarkon):

is this for a class you are taking?

myininaya (myininaya):

http://en.wikipedia.org/wiki/Computational_Diffie%E2%80%93Hellman_assumption here is a little more info on it --- i'm reading a paper

myininaya (myininaya):

i mean you don't have to look at that site if you aren't interested

OpenStudy (zarkon):

interesting

myininaya (myininaya):

yes it is

OpenStudy (anonymous):

zarkon, are you still here? I have another question that I could use some help on

OpenStudy (zarkon):

I might be here

myininaya (myininaya):

what kind of question?

myininaya (myininaya):

and i know i'm not zarkon lol

OpenStudy (zarkon):

tag team

OpenStudy (anonymous):

Prove the identity\[\arcsin (x-1/x+1) = 2\arctan \sqrt{x}-\pi/2\]

myininaya (myininaya):

oh i did that

myininaya (myininaya):

i can find it

myininaya (myininaya):

one sec

OpenStudy (zarkon):

yep I remember seeing this problem just a little while ago

myininaya (myininaya):

james did this the cal way i did it the trig way look at both and pick what you like http://openstudy.com/#/updates/4ec58ab6e4b03063799de754

OpenStudy (anonymous):

oo thank you

myininaya (myininaya):

do you have any questions on anything james or i did?

myininaya (myininaya):

i hope you like the trig way better because i think trig is the freaking bomb!

OpenStudy (anonymous):

I like the way you did it it was how I started doing the problem when I got stuck.

OpenStudy (anonymous):

calc way looks cool too though

myininaya (myininaya):

yeah its okay its like tapping out at a wrestling match but whatever you like

myininaya (myininaya):

i was jk

OpenStudy (anonymous):

the trig way just seems to work out so much better and easier, but you can get lost in all those trig formulas if you aren't careful lol

myininaya (myininaya):

right

myininaya (myininaya):

i bet my trig students wouldn't be able to figure this one out

OpenStudy (anonymous):

the arcsin and arctan make it tricky. I just recently learned that right triangle trick too.

myininaya (myininaya):

next semester i 'm going give to my trig and cal students as a project muhahaha however the evil laugh goes